# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v21

@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [439] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [440]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [441] which they later extended further [442]Epaliyana et al. [443] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [444] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [445] used a model based on fairseq [446] to improve machine translation between English and Sinhala. The second half of the study by Arafath [334] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [447] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [448]. Nguyen et al. [449] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [450]. Attigala and Weerasinghe [451] conducted an analysis on the effectiveness of ChatGPT [452] in translating Sinhala songs to English.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP106106106https://github.com/google-research/xtreme-up data set created by Ruder et al. [130] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [453] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [454] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [455] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [456]. A rule-based approach for the same was proposed by de Silva and Ahangama [457]. The study by Nanayakkara et al. [458] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [459, 460] proposed a Trie [461] data structure-based algorithm for word suggestion in Sinhala transliterations. An extended analysis of the same work was presented in a later work [462]. Amarasekara et al. [463] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [464] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [465] and Jayawardhana et al. [466] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [470, 471]. Then they further enhanced it with transliteration and byte pair encoding [472] and used synthetic training data to handle the rare word problem [473]. This project produced Si-Ta [474] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [475] worked on integrating bilingual lists. The attempts by Weerasinghe [476] and Sripirakas et al. [477] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [478] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [479] which they later extended with some quality improvements [480]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [481]Dilshani et al. [482] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [483] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [484] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [485]Pramodya et al. [486] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [487] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [488] uses pre-trained mBART [489] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [490] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [491] presented a comparison of SMT [436] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [287].
@@ 3.20 Spelling and Grammar
The open-source data driven approach proposed by Wasala et al. [498, 499] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [500] attempts to obtain synergy between two algorithms for the same purpose. These efforts [498, 500] were then extended by Subhagya et al. [501]. A rule-based Sinhala spell checker named SinSpell based on Hunspell107107107http://hunspell.github.io/ was introduced by Liyanapathirana et al. [502]. They have also made the tool available108108108http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [503] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [116] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [504]. The study by Udagedara et al. [505] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [506] to correct real-word errors in Sinhala text.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [67]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [516] where they used, the fasttext model trained by Facebook [109, 110, 111], on a RASA109109109https://rasa.com/ [517] chat bot. A Sinhala chat bot for train information was proposed by Harshani [518]. Similarly, the tool proposed by Chandrasena et al. [519] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [520] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [521] used off the shelf RASA NLU Engine [517] and Microsoft Bot Network [522] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [523]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [524] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [347] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
@@ 3.23 Sinhala Sign Language
Priyankara et al. [535] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [536] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [537], the approach by Dilakshan and Priyadarshana [538], as well as the approach by Peiris [539] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [540] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [541] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [542]. The study by Charuka et al. [139] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [543] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [544] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. Strides have been made in the domains of computer interpreting for written Sinhala [545] and animation of finger-spelled words and number signs [546].
Kumar et al. [547] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. Both the approaches by Perera and Kumarika [548] and Herath and Ishanka [549] integrated MediaPipe [550] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [548] only used LSTM while Herath and Ishanka [549] used LSTM, CNN, and CNN-LSTM. Rishan et al. [532] uses Leap Motion Controller110110110https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [551] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [562] and later by Smith [563] using an XGB classifier and a CRF model building on their previous work [564], which analysed such data. Shanmugalingam and Sumathipala [565] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [448]. They later used it for NMT in Sinhala-English Code-Mixed text [447]. As discussed in Section 3.10Rathnayake et al. [216] used adapter-based [217, 218, 219, 220, 221, 222] fine-tuning on XLM-R [223], for classifying code-mixed and code-switched Sinhala text. Hettigoda [226] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [246] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [228] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [523] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [388] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [566] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook.

### v22

@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [456] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [457]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [458] which they later extended further [459]Epaliyana et al. [460] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [461] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [462] used a model based on fairseq [463] to improve machine translation between English and Sinhala. The second half of the study by Arafath [348] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [464] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [465]. Nguyen et al. [466] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [467]. Attigala and Weerasinghe [468] conducted an analysis on the effectiveness of ChatGPT [469] in translating Sinhala songs to English. Utsa et al. [470] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [92] and reported better results than mBART [471].
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP114114114https://github.com/google-research/xtreme-up data set created by Ruder et al. [134] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [472] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [473] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [474] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [475]. A rule-based approach for the same was proposed by de Silva and Ahangama [476]. The study by Nanayakkara et al. [477] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [478, 479] proposed a Trie [480] data structure-based algorithm for word suggestion in Sinhala transliterations. An extended analysis of the same work was presented in a later work [481]. Amarasekara et al. [482] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [483] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [484] and Jayawardhana et al. [485] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
The work by Kirov et al. [486] used the Dakshina [146] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [299], ByT5 [487], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [488]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub115115115https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [489] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [493, 494]. Then they further enhanced it with transliteration and byte pair encoding [495] and used synthetic training data to handle the rare word problem [496]. This project produced Si-Ta [497] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [498] worked on integrating bilingual lists. The attempts by Weerasinghe [499] and Sripirakas et al. [500] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [501] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [502] which they later extended with some quality improvements [503]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [504]Dilshani et al. [505] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [506] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [507] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [508]Pramodya et al. [509] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [510] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [511] uses pre-trained mBART [471] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [512] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [513] presented a comparison of SMT [453] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [299]. The work by Su et al. [514] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [454] and NLLB [127] reports that Houlsby adapter [224], with a 33.34 SacreBLEU [515] score, to be the best PEFT method. However, they also note that the Pfeiffer adapter [226] runs the fastest at 52.59 hours with a reasonable score of 31.24. (Comparatively, the Houlsby adapter runs for 78.65 hours).
@@ 3.20 Spelling and Grammar
The open-source data driven approach proposed by Wasala et al. [522, 523] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [524] attempts to obtain synergy between two algorithms for the same purpose. These efforts [522, 524] were then extended by Subhagya et al. [525]. A rule-based Sinhala spell checker named SinSpell based on Hunspell116116116http://hunspell.github.io/ was introduced by Liyanapathirana et al. [526]. They have also made the tool available117117117http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [527] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [120] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [528]. The study by Udagedara et al. [529] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [530] to correct real-word errors in Sinhala text.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [70]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [540] where they used, the fasttext model trained by Facebook [113, 114, 115], on a RASA118118118https://rasa.com/ [541] chat bot. A Sinhala chat bot for train information was proposed by Harshani [542]. Similarly, the tool proposed by Chandrasena et al. [543] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [544] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [545] used off the shelf RASA NLU Engine [541] and Microsoft Bot Network [546] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [547]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [548] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [362] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
@@ 3.23 Sinhala Sign Language
Priyankara et al. [559] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [560] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [561], the approach by Dilakshan and Priyadarshana [562], as well as the approach by Peiris [563] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [564] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [565] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [566]. The study by Charuka et al. [143] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [567] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [568] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. The study by Fernando et al. [569] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [570] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [571] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [572] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [573]. The work by Perera [574] uses CNN for Sri Lanka Sign Language recognition. Strides have been made in the domains of computer interpreting for written Sinhala [575] and animation of finger-spelled words and number signs [576]. The work by Idushan et al. [577] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
Kumar et al. [578] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. Both the approaches by Perera and Kumarika [579] and Herath and Ishanka [580] integrated MediaPipe [581] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [579] only used LSTM while Herath and Ishanka [580] used LSTM, CNN, and CNN-LSTM. Rishan et al. [556] uses Leap Motion Controller119119119https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [582] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [593] and later by Smith [594] using an XGB classifier and a CRF model building on their previous work [595], which analysed such data. Shanmugalingam and Sumathipala [596] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [465]. They later used it for NMT in Sinhala-English Code-Mixed text [464]. As discussed in Section 3.10Rathnayake et al. [223] used adapter-based [224, 225, 226, 227, 228, 229] fine-tuning on XLM-R [230], for classifying code-mixed and code-switched Sinhala text. Hettigoda [233] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [253] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [235] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [547] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [404] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [597] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by uda [598] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection.
3.27 Language Acquisition and Evaluation
Sandathara et al. [599] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [600] assists children to learn to speak without stuttering. Nethmi et al. [601] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [602] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [603] uses SVM and LSTm to identify phonological disorders from Sinhala speech.
3.28 Embeddings and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [604]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [101] has published two Sinhala BERT models: SinBERT-large120120120https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small121121121https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo122122122https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [605] on the OSCAR [144, 145] dataset.
Üstün et al. [606] introduced Aya, a massively multilingual generative language model123123123https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [148].
Wickramasinghe and de Silva [138] used Sinhala-English dictionary data [137] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub124124124https://bit.ly/3t3SKu7.